# Optimal Threshold for Fracton Codes and Nearly Saturated Code Capacity in Three Dimensions

### v1

@@ IV.1 Random Spin Models With Subsystem Symmetry
In the checkerboard code, each qubit is shared by four stabilizers; using the method outlined in section III.2, the error model can be mapped into a random tetrahedral Ising model [Canossa24, Vijay15]. For both the $X$ and $Z$ error model, each qubit $i$ lies at the corner of 4 stabilizers of the checkerboard code: in the effective spin model, we will label the set of Ising degrees of freedom associated to the stabilizer sharing said qubit $i$ as $\sigma_{j}$$\sigma_{k}$$\sigma_{l}$$\sigma_{m}$$\mathcal{N}_{X}(i)=\mathcal{N}_{Z}(i)$.
@@
$\mathcal{Z}(\eta,T)=\sum_{\{\sigma\}}\text{exp}\left[\beta\sum_{i\in Q}\eta_{i}\,\,\sigma_{j}\sigma_{k}\sigma_{l}\sigma_{m}\right]$ (17)
@@
Figure 2: Illustration of the statistical mapping from the checkerboard code (left) to the tetrahedral Ising model on a face-centered cubic lattice (right) derived from [Canossa24]. Blue and cyan circles, located at the vertices and face centers of the FCC lattice, represent the Ising spins associated with each stabilizer, constructing the dual lattice of the checkerboard code. The action of a single qubit error is mapped into a four-body interaction with strength $\eta_{\mathbf{v}}^{\pm}$; these are defined for each lattice site $\mathbf{v}$ in the FCC lattice and are represented by a red (green) tetrahedron placed in each shaded cube (blank cube) of the lattice on the right.
@@ IV.2 Disorder Simulations
The tetrahedral Ising model with no disorder is characterized by a very strong first-order phase transition [Canossa24]. While quenched disorder reduces the amount of latent heat required to break the phase boundaries, the features of the first-order transition are still relatively sharp near the threshold (see Fig. 3): this leads to very slow equilibration which makes the simulation of each disorder configuration of the error model severely expensive from a computational perspective. When paired to the fact that one needs to study the average of data from several hundreds of disorder iterations in order to study the disorder average, it is necessary to optimize the individual simulations in order to minimize their equilibration time and overall runtime. With this in mind, each disorder realization has been simulated with parallel tempering in conjunction with Metropolis-Hastings and overrelaxation updates in order to speed up the equilibration process [Katzgraber06, Janke08]. The temperature distribution is tailored for all disorder simulations at a given disorder value $p$ and system size $L$ in order to maximize the mixing of replicas between ordered and disordered phase in the average case.
The disorder averages are obtained through a large number of configurations ($N_{d}\sim 500$), and their statistical uncertainties are derived through the Jackknife resampling method [Janke08]. The specific number of temperatures and disorder iterations used for each disorder average is reported in Table 2.
Equilibration of each individual disorder iteration is inferred by combining various analyses. The timeseries of the energy is divided into bins with intervals $2^{\tau},2^{\tau+1}-1]$, where $\tau\in\mathbb{N}$ labels a specific bin. After averaging the values stored in each bin, the system can be considered equilibrated when the last values of the last three bins are within error bars of each other [Bombin12].
@@
In the few cases where this was not sufficient, the energy histograms were tested through a reweighting method [Janke08]: after collecting the energy histogram for each inverse temperature $\beta$, the energy histogram at a new temperature $\beta^{\prime}$ can be derived by applying to each bin at energy $E$ a weight $e^{-(\beta^{\prime}-\beta)E}$. The simulation is considered equilibrated if, for most temperatures near the phase transition, the energy histograms agree within sufficient statistical uncertainty after appropriate reweighting. This ensures that, while the disorder average at certain individual temperatures might still have gotten stuck in one of the two phases, the majority of the replicas have consistently captured the expected behaviour of the system at their respective temperatures.
@@
After taking into account the optimization of the temperature distribution, the system still required a remarkable amount of Monte Carlo sweeps to reach equilibration. The simulations were run on the LRZ KCS cluster [LRZ_Linux_Cluster] and used over $7\times 10^{6}$ CPU hours in total, with an average of $\approx 10^{6}$ CPU hours of simulation time per $(L,p)$ dataset. Due to the cumbersome amount of computation time each point requires, computation was restricted to a limited set of $(L,p)$ points. The system sizes simulated lie in the range of $L\in[10,14]$, as lower system sizes were affected by excessively large finite-size effects which make them irrelevant for scaling purposes whereas higher system sizes were computationally unapproachable due to the exponential growth of the simulation times required. After a preliminary round of simulations used to infer a rough position of the error threshold, the range of relevant disorder values was originally set to be $p\in[0.105,0.110]$. To achieve an estimate of the threshold value with precision comparable to that of previous works [Katzgraber09, Kubica18, Song22], it was sufficient to simulate only 3 datapoints: the lower bound, an upper bound, and a conservative estimate of the possible location of the threshold, which was set at $p^{\text{mid}}=0.107$. Simulating additional intermediate disorder values would only increase computational cost without improving the statistical robustness of the threshold estimate: for this reason, the reported simulation set represents the minimal dataset that captures all qualitative changes relevant for locating the tricritical region while keeping the computational effort within feasible limits.
@@ IV.3 Numerical Results
We simulated the random TIM using large-scale parallel tempering Monte Carlo simulations, averaging the results of enough disorder configurations in order to obtain the disorder averaged features of the model. Each configuration simulated for a given value of the disorder parameter $p$ is obtained by starting from a fully FM set of couplings $(\eta_{\mathbf{v}}^{\pm}=+1\,\,\forall\,\,\mathbf{v})$ and then proposing a random flip to each coupling with probability $p$. This way, the average ratio between number of AFM couplings and the total number of couplings equals $p$.
Accompanying the phenomenon of subsystem symmetry breaking [Canossa23, Canossa24], the model shows strong features of a first-order phase transition even in proximity of the error threshold. This property, coupled with the restricted set of lattice sizes accessible, requires us to carefully analyze the subtle changes in the thermodynamic features of the model between the system sizes available. To infer whether the system shows a transition at a given disorder value, we studied the behavior of the energy histogram
$P(E)=\delta\left(E-E^{\prime}\right),$ (19)
and the second-moment correlation length
$\xi_{L}=\frac{1}{2\sin\left(|\textbf{k}_{\text{min}}|\right)}\left(\frac{\tilde{G}(\mathbf{0})}{\tilde{G}(\textbf{k}_{\text{min}})}-1\right)^{1/2}.$ (20)
Here, $\tilde{G}(\textbf{k}):=\sum_{\textbf{r}}G(\textbf{r})e^{-i\textbf{k}\cdot\textbf{r}}$ is the Fourier transform of the spatial correlator $G(\textbf{r})$ and $\textbf{k}_{\text{min}}$ is the smallest non-zero wave vector in reciprocal space [Janke08]. The spatial correlator of the model is defined as [Canossa24]
$G(r)=\frac{1}{N}\sum_{\bf v}\left\langle\sigma_{\bf v}\sigma_{{\bf v}+\hat{x}+\hat{y}}\sigma_{{\bf v}+\hat{y}+r\hat{z}}\sigma_{{\bf v}+\hat{x}+r\hat{z}}\right\rangle$ (21)
@@
In order to prove whether the checkerboard code saturates the QGV bound, the search for the error threshold was done at first for disorder values around $p=0.11$, the expected critical point for self-dual models with quenched disorder [Takeda05, Nishimori07, Song22]. At $p=0.110$ and above, the double peak behavior in the energy histograms is still visible, but tends to disappear at larger system sizes. This, combined with the lack of crossing in the correlation length (which rules out the possible presence of a continuous phase transition), confirms that the system is indeed in a crossover regime and that the double-peak behavior observed is caused by finite-size effects and will disappear in the thermodynamic limit. This serves as a characteristic example of the behavior of the system in the disordered regime and a reference against which to compare observations at lower error rates, especially due to the fact that the decisive signatures lie in the ways in which the strong finite-size effects evolve as the system size increases.
@@
At disorder value $p=0.105$ and below, the two peaks tend sharpen and the gap between them deepens at the system size increases, indicating the presence of a first-order phase transition. Given the first-order nature shown by the energy histograms, a crossing in the correlation length is not necessarily expected. Nevertheless, the observation of a crossing between the larger system sizes is consistent with the relaxation of finite-size effects and rules out the eventuality of a crossover regime.
@@
The behavior shown at $p=0.107$ was similar, although much less accentuated, as the minimal point in the gap of the energy histograms lowers only slightly with system sizes; considering the presence of a crossing in the correlation lengths between the two curves at larger system sizes, we argue that the system is still in the decodable phase, but is in close proximity to the threshold.
@@
These observations allow us to estimate the tricritical point. Since we can place it within the interval $(0.107,0.110)$, the most conservative claim that does not exclude any value in this range is to set it at $p_{th}\simeq 0.108(2)$. The size of the error bar is comparable to that in previous numerical threshold estimates (see Table 1) and is sufficient to cover the full interval of uncertainty while remaining deliberately conservative.
@@
This threshold represents the highest optimal error threshold among any $3D$ topological codes. Moreover, it gives to $H(p_{th}^{X})+H(p_{th}^{Z})\approx 0.987(8)$, which nearly saturates the QGV bound and also confirms the applicability of the generalized duality for the checkerboard code.
@@
Figure 5: (a) Illustration of the stabilizer structure of the Haah’s code. Two sets qubits are situated at the vertices of a cubic lattice. The $X$- and $Z$- type stabilizers $A_{C},B_{C}$ are defined for each unit cell $C$ as a product of Pauli operators acting on a subset of the 16 qubits residing at the vertices of the cube. (b) A single-qubit $X$ (or $Z$) error excites 4 cubic stabilizer operators: since the structures are equivalent for $X$ and $Z$ stabilizers up to a global rotation, both independent noise models can be represented by random the Fractal Ising Model Eq. (V).
@@ V Considerations about Haah’s code
Haah’s cubic code is a seminal fracton code, notable for having no string-like logical operators and for being one of the first constructions to reveal fracton order in a quantum error-correction context [Haah11]. The code is defined on a cubic lattice with two qubits per site, and its stabilizers follow the general structure introduced in Eq. (1): each cube hosts an $X$- and $Z$-type 8-body stabilizer. Identifying each cube $C$ with its $(0,0,0)$ vertex $\bf v$, and labelling the two families of qubits as $\{1,2\}$, the stabilizers take the form
$\hat{A}_{C}$ $=\hat{\sigma}^{X,2}_{\mathbf{v}+\hat{x}}\,\hat{\sigma}^{X,2}_{\mathbf{v}+\hat{x}}\,\hat{\sigma}^{X,1}_{\mathbf{v}+\hat{x}+\hat{y}}\,\hat{\sigma}^{X,2}_{\mathbf{v}+\hat{z}}\,\,\,\,\,\,\,\hat{\sigma}^{X,1}_{\mathbf{v}+\hat{z}+\hat{x}}\,\hat{\sigma}^{X,1}_{\mathbf{v}+\hat{z}+\hat{y}}\,\hat{\sigma}^{X,1}_{\mathbf{v}+\hat{x}+\hat{y}+\hat{z}}\,\hat{\sigma}^{X,2}_{\mathbf{v}+\hat{x}+\hat{y}+\hat{z}}$
$\hat{B}_{C}$ $=\hat{\sigma}^{Z,1}_{\mathbf{v}}\,\,\,\hat{\sigma}^{Z,2}_{\mathbf{v}}\,\,\,\hat{\sigma}^{Z,2}_{\mathbf{v}+\hat{x}}\,\,\,\,\,\,\hat{\sigma}^{Z,2}_{\mathbf{v}+\hat{x}+\hat{y}}\,\hat{\sigma}^{Z,1}_{\mathbf{v}+\hat{x}+\hat{y}}\,\hat{\sigma}^{Z,2}_{\mathbf{v}+\hat{z}}\,\,\,\,\,\,\,\hat{\sigma}^{Z,1}_{\mathbf{v}+\hat{z}+\hat{x}}\,\,\,\,\,\,\hat{\sigma}^{Z,1}_{\mathbf{v}+\hat{z}+\hat{y}}$
The model is characterized by a subextensive number of fractal symmetries [Haah13]: as a consequence of these $\mathbb{Z}_{2}$ fractal conservation law, a single-qubit error creates a cluster of four fracton excitations which cannot be separated by any local operator. These defects can only move collectively along the underlying fractal pattern and are therefore individually immobile.
@@
with integers $n,m\geq 1$ [Canossa24]. This strongly affects the behavior of the system, leading to different finite-size effects for the different system size sequences. Even for the simplest size sequence $L=2^{n}$, the accessible lattice sizes are restricted to $L=4,8$ which are insufficient to locate the phase transition due to strong finite-size effects. Larger systems ($L\geq 16$) are computationally prohibitive, with preliminary runtime estimates approaching $\approx 1.5\times 10^{7}$ additional CPU hours of computation per $(L=16,p)$ dataset.

### v2

@@ IV.1 Random Spin Models With Subsystem Symmetry
In the checkerboard code, each qubit is shared by four stabilizers; using the method outlined in section III.2, the error model can be mapped into a random Tetrahedral Ising model [15, 57]. For both the $X$ and $Z$ error model, each qubit $i$ lies at the corner of 4 stabilizers of the checkerboard code: in the effective spin model, we will label the set of Ising degrees of freedom associated to the stabilizer sharing said qubit $i$ as $\sigma_{j}$$\sigma_{k}$$\sigma_{l}$$\sigma_{m}$$\mathcal{N}_{X}(i)=\mathcal{N}_{Z}(i)$.
@@
$\mathcal{Z}(\eta,T)=\sum_{\{\sigma\}}\text{exp}\left[\beta\sum_{i\in Q}\eta_{i}\,\,\sigma_{j}\sigma_{k}\sigma_{l}\sigma_{m}\right].$ (17)
@@
Figure 2: Illustration of the statistical mapping from the checkerboard code (left) to the Tetrahedral Ising model on a face-centered cubic lattice (right) derived from [15]. Blue and cyan circles, located at the vertices and face centers of the FCC lattice, represent the Ising spins associated with each stabilizer, constructing the dual lattice of the checkerboard code. The action of a single qubit error is mapped into a four-body interaction with strength $\eta_{\mathbf{v}}^{\pm}$; these are defined for each lattice site $\mathbf{v}$ in the FCC lattice and are represented by a red (green) tetrahedron placed in each shaded cube (blank cube) of the lattice on the right.
@@ IV.2 Disorder Simulations
The Tetrahedral Ising model with no disorder is characterized by a very strong first-order phase transition [15]. While quenched disorder reduces the amount of latent heat required to break the phase boundaries, the features of the first-order transition are still relatively sharp near the threshold (see Fig. 3): this leads to very slow equilibration which makes the simulation of each disorder configuration of the error model severely expensive from a computational perspective. When paired to the fact that one needs to study the average of data from several hundreds of disorder iterations in order to study the disorder average, it is necessary to optimize the individual simulations in order to minimize their equilibration time and overall runtime. With this in mind, each disorder realization has been simulated with parallel tempering in conjunction with Metropolis-Hastings and overrelaxation updates in order to speed up the equilibration process [32, 31]. The temperature distribution is tailored for all disorder simulations at a given disorder value $p$ and system size $L$ in order to maximize the mixing of replicas between ordered and disordered phase in the average case.
The disorder averages are obtained through a large number of configurations ($N_{d}\sim 500$), and their statistical uncertainties are derived through the Jackknife resampling method [31]. The specific number of temperatures and disorder iterations used for each disorder average is reported in Table 2.
Equilibration of each individual disorder iteration is inferred by combining various analyses. The timeseries of the energy is divided into bins with intervals $2^{\tau},2^{\tau+1}-1]$, where $\tau\in\mathbb{N}$ labels a specific bin. After averaging the values stored in each bin, the system can be considered equilibrated when the last values of the last three bins are within error bars of each other [5].
@@
In the few cases where this was not sufficient, the energy histograms were tested through a reweighting method [31]: after collecting the energy histogram for each inverse temperature $\beta$, the energy histogram at a new temperature $\beta^{\prime}$ can be derived by applying to each bin at total energy $E$ a weight $e^{-(\beta^{\prime}-\beta)E}$. The simulation is considered equilibrated if, for most temperatures near the phase transition, the energy histograms agree within sufficient statistical uncertainty after appropriate reweighting. This ensures that, while the disorder average at certain individual temperatures might still have gotten stuck in one of the two phases, the majority of the replicas have consistently captured the expected behaviour of the system at their respective temperatures.
@@
After taking into account the optimization of the temperature distribution, the system still required a remarkable amount of Monte Carlo sweeps to reach equilibration. The simulations were run on the LRZ KCS cluster [38] and used over $7\times 10^{6}$ CPU hours in total, with an average of $\approx 10^{6}$ CPU hours of simulation time per $(L,p)$ dataset. Due to the cumbersome amount of computation time each point requires, computation was restricted to a limited set of $(L,p)$ points. The system sizes simulated lie in the range of $L\in[10,14]$, as lower system sizes were affected by excessively large finite-size effects which make them irrelevant for scaling purposes whereas higher system sizes were computationally unapproachable due to the exponential growth of the simulation times required. After a preliminary round of simulations used to infer a rough position of the error threshold, the range of relevant disorder values was originally set to be $p\in[0.105,0.110]$. To achieve an estimate of the threshold value with precision comparable to that of previous works [33, 36, 53], it was sufficient to simulate only 3 datapoints: the lower bound, an upper bound, and a conservative estimate of the possible location of the threshold, which was set at $p^{\text{mid}}=0.107$. Simulating additional intermediate disorder values would only increase computational cost without improving the statistical robustness of the threshold estimate: for this reason, the reported simulation set represents the minimal dataset that captures all qualitative changes relevant for locating the tricritical region while keeping the computational effort within feasible limits.
@@ IV.3 Numerical Results
We simulated the random Tetrahedral Ising model using large-scale parallel tempering Monte Carlo simulations, averaging the results of enough disorder configurations in order to obtain the disorder averaged features of the model. Each configuration simulated for a given value of the disorder parameter $p$ is obtained by starting from a fully FM set of couplings $(\eta_{\mathbf{v}}^{\pm}=+1\,\,\forall\,\,\mathbf{v})$ and then proposing a random flip to each coupling with probability $p$. This way, the average ratio between number of AFM couplings and the total number of couplings equals $p$.
Accompanying the phenomenon of subsystem symmetry breaking [14, 15], the model shows strong features of a first-order phase transition even in proximity of the error threshold. This property, coupled with the restricted set of lattice sizes accessible, requires us to carefully analyze the subtle changes in the thermodynamic features of the model between the system sizes available. To infer whether the system shows a transition at a given disorder value, we studied the behavior of the energy histogram
$P(e)=\langle\delta\left(e-e^{\prime}\right)\rangle,$ (19)
where $e^{\prime}=E^{\prime}/N$ is the energy density for a given configuration of a system of size $N=L^{3}/2$, and the second-moment correlation length
$\xi_{L}=\frac{1}{2\sin\left(|\textbf{k}_{\text{min}}/2|\right)}\left(\frac{\tilde{G}(\mathbf{0})}{\tilde{G}(\textbf{k}_{\text{min}})}-1\right)^{1/2}.$ (20)
Here, $\tilde{G}(\textbf{k}):=\sum_{\textbf{r}}G(\textbf{r})e^{-i\textbf{k}\cdot\textbf{r}}$ is the Fourier transform of the spatial correlator $G(\textbf{r})$ and $\textbf{k}_{\text{min}}$ is the smallest non-zero wave vector in reciprocal space [31]. The spatial correlator of the model is defined as [15]
$G(r)=\frac{1}{N}\sum_{\bf v}\left\langle\sigma_{\bf v}\sigma_{{\bf v}+\hat{x}+\hat{y}}\sigma_{{\bf v}+\hat{y}+r\hat{z}}\sigma_{{\bf v}+\hat{x}+r\hat{z}}\right\rangle$. (21)
@@
In order to prove whether the checkerboard code saturates the threshold bound, the search for the error threshold was done at first for disorder values around $p=0.11$, the expected critical point for self-dual models with quenched disorder [55, 43, 53]. At $p=0.110$ and above, the double peak behavior in the energy histograms is still visible, but tends to disappear at larger system sizes. This, combined with the lack of crossing in the correlation length (which rules out the possible presence of a continuous phase transition), confirms that the system is indeed in a crossover regime and that the double-peak behavior observed is caused by finite-size effects and will disappear in the thermodynamic limit. This serves as a characteristic example of the behavior of the system in the disordered regime and a reference against which to compare observations at lower error rates, especially due to the fact that the decisive signatures lie in the ways in which the strong finite-size effects evolve as the system size increases.
@@
At disorder value $p=0.105$ and below, the two peaks tend to sharpen and the gap between them deepens as the system size increases, indicating the presence of a first-order phase transition. Given the first-order nature shown by the energy histograms, a crossing in the correlation length is not necessarily expected. Nevertheless, the observation of a crossing between the larger system sizes is consistent with the relaxation of finite-size effects and rules out the eventuality of a crossover regime.
@@
The behavior shown at $p=0.107$ was similar, although much less accentuated, as the minimal point in the gap of the energy histograms lowers only slightly with system sizes; considering the presence of a crossing in the correlation lengths between the two curves at larger system sizes, we argue that the system is still in the decodable phase, but is in very close proximity to the threshold.
@@
These observations allow us to estimate the tricritical point. Since we can place it within the interval $[0.105,0.110)$, the most conservative claim that does not exclude any value in this range is to set it at $p_{th}\simeq 0.107(3)$. The size of the error bar is comparable to that in previous numerical threshold estimates (see Table 1) and is sufficient to cover the full interval of uncertainty while remaining deliberately conservative.
@@
This threshold represents the highest optimal error threshold among any $3D$ topological codes. Moreover, it gives to $H(p_{th}^{X})+H(p_{th}^{Z})\approx 0.98(2)$, which nearly saturates the error threshold bound and also confirms the applicability of the generalized duality for the checkerboard code.
@@
Figure 5: (a) Illustration of the stabilizer structure of the Haah’s code. Two sets qubits are situated at the vertices of a cubic lattice. The $X$- and $Z$- type stabilizers $A_{C},B_{C}$ are defined for each unit cell $C$ as a product of Pauli operators acting on a subset of the 16 qubits residing at the vertices of the cube. (b) A single-qubit $X$ (or $Z$) error excites 4 cubic stabilizer operators: since the structures are equivalent for $X$ and $Z$ stabilizers up to a global rotation, both independent noise models can be represented by the random Fractal Ising Model Eq. (V).
@@ V Considerations about Haah’s code
Haah’s cubic code is a seminal fracton code, notable for having no string-like logical operators and for being one of the first constructions to reveal fracton order in a quantum error-correction context [25]. The code is defined on a cubic lattice with two qubits per site, and its stabilizers follow the general structure introduced in Eq. (1): each cube hosts an $X$- and $Z$-type 8-body stabilizer. Identifying each cube $C$ with its $(0,0,0)$ vertex $\bf v$, and labelling the two families of qubits as $\{1,2\}$, the stabilizers take the form
$\hat{A}_{C}$ $=\hat{\sigma}^{X,2}_{\mathbf{v}+\hat{x}}\,\hat{\sigma}^{X,2}_{\mathbf{v}+\hat{x}}\,\hat{\sigma}^{X,1}_{\mathbf{v}+\hat{x}+\hat{y}}\,\hat{\sigma}^{X,2}_{\mathbf{v}+\hat{z}}\,\,\,\,\hat{\sigma}^{X,1}_{\mathbf{v}+\hat{z}+\hat{x}}\,\hat{\sigma}^{X,1}_{\mathbf{v}+\hat{z}+\hat{y}}\,\hat{\sigma}^{X,1}_{\mathbf{v}+\hat{x}+\hat{y}+\hat{z}}\,\hat{\sigma}^{X,2}_{\mathbf{v}+\hat{x}+\hat{y}+\hat{z}},$
$\hat{B}_{C}$ $=\hat{\sigma}^{Z,1}_{\mathbf{v}}\,\,\,\hat{\sigma}^{Z,2}_{\mathbf{v}}\,\,\,\hat{\sigma}^{Z,2}_{\mathbf{v}+\hat{x}}\,\,\,\,\hat{\sigma}^{Z,2}_{\mathbf{v}+\hat{x}+\hat{y}}\,\hat{\sigma}^{Z,1}_{\mathbf{v}+\hat{x}+\hat{y}}\,\hat{\sigma}^{Z,2}_{\mathbf{v}+\hat{z}}\,\,\,\,\hat{\sigma}^{Z,1}_{\mathbf{v}+\hat{z}+\hat{x}}\,\,\,\,\,\,\hat{\sigma}^{Z,1}_{\mathbf{v}+\hat{z}+\hat{y}}.$
The model is characterized by a subextensive number of fractal symmetries [26]: as a consequence of these $\mathbb{Z}_{2}$ fractal conservation law, a single-qubit error creates a cluster of four fracton excitations which cannot be separated by any local operator. These defects can only move collectively along the underlying fractal pattern and are therefore individually immobile.
@@
with integers $n,m\geq 1$ [15]. This strongly affects the behavior of the system, leading to different finite-size effects for the different system size sequences. Even for the simplest size sequence $L=2^{n}$, the accessible lattice sizes are restricted to $L=4,8$ which are insufficient to locate the phase transition due to strong finite-size effects. Larger systems ($L\geq 16$) are computationally prohibitive, with preliminary runtime estimates approaching $\approx 1.5\times 10^{7}$ additional CPU hours of computation per $(L=16,p)$ dataset.